Find the impulse response function. y +8y + 16y = g(t)
Answer #1

y+8yt 16y=gw) g) = S(4) Limpulse response) Y(0)=0, gion=1 Homogeneous solen y+8yt by=0; y=emx. mt 8m +16=0 (m+35=0 m=74